# Detecting Neurocognitive Disorders through Analyses of Topic Evolution and Cross-modal Consistency in Visual-Stimulated Narratives

### v1

@@ Abstract
Early detection of neurocognitive disorders (NCDs) is crucial for timely intervention and disease management. Since language abilities are affected by NCDs at a very early stage, speech analysis emerges as a non-intrusive, and scalable method for screening NCDs, especially through narrative tasks that hold significant potential for evaluating cognitive abilities. Traditional narrative analysis often focuses on local indicators in microstructure, such as word choice and syntax. While these features provide insights into language production abilities, they often fail to capture global narrative patterns, or microstructures. Macrostructures include coherence, thematic organization, and logical progressions, reflecting essential cognitive skills potentially critical for recognizing NCDs. Addressing this gap, we propose to investigate specific cognitive and linguistic challenges by analyzing topical shifts, temporal dynamics, and the coherence of narratives over time, aiming to reveal cognitive deficits by identifying narrative impairments, and exploring their impact on communication and cognition. The investigation is based on the CU-MARVEL Rabbit Story corpus, which comprises recordings of a story-telling task from 758 older adults. We developed two approaches: the Dynamic Topic Models (DTM)-based temporal analysis to examine the evolution of topics over time, and the Text-Image Temporal Alignment Network (TITAN) to evaluate the coherence between spoken narratives and visual stimuli. DTM-based approach validated the effectiveness of dynamic topic consistency as a macrostructural metric (F1=0.61, AUC=0.78). The TITAN approach achieved the highest performance (F1=0.72, AUC=0.81), surpassing established microstructural and macrostructural feature sets. Cross-comparison and regression tasks further demonstrated the effectiveness of proposed dynamic macrostructural modeling approaches for NCD detection.
@@ I Introduction
Neurocognitive disorders (NCDs)[1], also known as dementia, pose significant challenges to individuals and healthcare systems in our aging society. Early detection of these conditions is critical, as it can allow for timely intervention, potentially slowing disease progression and improving the quality of life for patients[2]. Traditional clinical approaches such as neuroimaging [3, 4], retinal photographs [5], electroencephalography (EEG) [6], cerebrospinal fluid analysis [7], and in-person neuropsychological assessments [8] have been widely investigated. However, these approaches can be resource-intensive, invasive, and costly [9, 10]. Consequently, there is a growing emphasis on developing non-invasive, cost-effective, and scalable methods for early NCD detection.
One promising approach is the automatic analysis of speech and language, as communicative impairments are often among the earliest symptoms of cognitive decline [11]. These impairments may include temporal changes in spontaneous speech [12], word finding and retrieval difficulties [13], verbal dysfluencies [14], reduction in discourse-level processing abilities [15, 16]. In light of this, various spoken cognitive tests have been designed to discover spoken language correlates for NCD detection. Such changes in speech patterns and language use provide valuable insights into underlying cognitive deficits and serve as potential biomarkers for early neurocognitive disorder (NCD) detection. In light of this, various spoken cognitive tests have been developed to uncover spoken language correlates for NCD detection. Traditional tests mostly focus on specific, discrete language functions, such as counting[17] and naming[18], where participants are required to perform highly structured and narrowly defined tasks. While these tests are useful for evaluating isolated aspects of language ability, they are limited in their scope and do not allow participants to fully demonstrate their comprehensive language capabilities. Moreover, the constrained nature of these tasks may fail to capture the dynamic, context-driven aspects of language use that are critical for assessing real-world communication skills. This highlights the need for holistic approaches to language assessment in naturalistic settings. More open-ended tasks, such as the visual-stimulated narrative (VSN) task, provide a versatile method for evaluating spontaneous language in individuals for assessment of cognitive functioning.
The VSN task involves describing a single image or telling a story based on a sequence of images, capturing an array of narrative abilities [19, 20, 21]. These narratives are deeply tied to neurocognitive domains through processes like describing, sequencing, and integrating elements. Producing a coherent narrative requires multiple cognitive functions, including attention, visual comprehension, language fluency, executive planning, memory, and social cognition [22, 23, 24, 25, 26, 27, 28]. This high cognitive demand poses challenges for individuals with NCDs, who often have deficits in these areas  [29]. Consequently, VSNs can potentially be used as a critical tool for early NCD detection [30, 11].
Approaches to narrative analysis can be categorized into two complementary aspects  [31]. First, the mainstream approach is microstructural narrative analysis, which examines linguistic (e.g., lexicon, semantics, syntax) [32, 33] and acoustic (e.g., prosody, phonology) [34, 35] features. In contrast, macrostructural analysis evaluates the overall structure and organization of storyline, emphasizing coherence (logical flow of concepts) and consistency (alignment with visual stimuli) [36]. Clinical research indicates that individuals with NCDs often exhibit greater impairment in macrostructure than microstructure, given that the microstructure provides more insights into their language abilities (perception and production), while the macrostructure offers more information on the cognitive abilities (attention, memory, organization, etc.) [31]. To date, macrostructural analysis remains underexplored, particularly within the context of NCD narratives.
@@
Macrostructural indicators like coherence and consistency are challenging to quantify due to their abstract and conceptual nature. These indicators involve assessing how well narrative elements align with visual stimuli and how logically topics are interconnected and evolved along story progression, requiring complex cognitive processing. Traditional approaches to coherence rely on manual assessments, which are subjective and labor-intensive [37]. Recent advances in natural language processing (NLP) have introduced automated methods, such as averaging embedding similarities or using large language models (LLMs) as evaluators [38, 39]. However, these methods often fall short for NCD narratives. Embedding-based similarity can misinterpret frequent repetitions or fillers as better cohesion and ignore dynamic evolutions of narrative topics over time. LLM-generated scores lack interpretability, undermining reliability in clinical applications.
@@
For consistency between narratives and visual stimuli, recent studies have used text-visual pretrained models to identify the highest relevance scores between image regions and sentences [40] or pattern-matching with fixed references [41, 42]. However, these methods have limitations: local text-image relevance fails to capture global consistency, and pattern-matching is restricted to fixed-reference tasks, making it unsuitable for creative tasks like storytelling. Therefore, tailored and automated approaches are needed for macrostructural analysis of NCD narratives.
@@
This work proposed to investigate specific cognitive and linguistic challenges by analyzing topical shifts, their temporal dynamics, and the coherence of narratives over time. We presented two novel approaches to macrostructural narrative analysis: Dynamic Topic Models (DTM) for temporal analysis of topical evolution; and Text-Image Temporal Alignment Model (TITAN) for cross-modal (i.e. stimulus images and spoken narratives) alignment analysis. This approach aims to reveal cognitive deficits by identifying narrative impairments associated with NCDs, and to provide a deeper understanding of their impact on communication and cognition. The investigations are conducted using the Rabbit Story task (a VSN task) from the CU-MARVEL corpus (a Cantonese corpus), as well as the single-image description task in the ADReSS corpus (an English corpus). Our experiments validated the effectiveness of dynamic topic consistency as a macrostructure metric (F1=0.61, AUC=0.78). The proposed TITAN model achieved the highest performance (F1=0.72, AUC=0.81) in classification tasks, surpassing established microstructure and macrostructure feature sets. Moreover, experiments of cross-comparison and regression tasks further demonstrated the effectiveness of proposed dynamic macrostructure modeling approaches for NCD detection.
@@
The paper is organized as follows: section II reviews related works, section III introduces corpora used in this study, section IV introduce the proposed approaches, section V presents our experimental settings and results, section VI analysis and discusses experimental results, and section VII concludes this work and prospects future works.
@@ II Related Work
Extensive research has been conducted for NCD detection based on spoken language, leveraging various methodologies to identify biomarkers and assess cognitive impairments. This section reviews pertinent literature encompassing both microstructural and macrostructural narrative analysis, collectively laying the groundwork for enhanced narrative analyses. The term “microstructure” refers to fine-grained elements of narrative production, including lexical, syntax, semantics, phonetics, or acoustics, and other smaller linguistic units. In contrast, “macrostructure” refers to higher-level aspects of narrative construction, such as coherence, cohesion, thematic development, and the overall organization of narratives. Both microstructural and macrostructural analysis are crucial for a comprehensive understanding of cognitive deficits reflected in NCD narratives. However, existing studies often face limitations in capturing the full spectrum of narrative features, particularly the dynamic aspects of narrative construction. By identifying these limitations, this review highlights the gaps the proposed methods aim to fill, specifically by incorporating temporal dynamics and macrostructural consistency, thus advancing the narrative analysis of NCDs.
@@ II-A Microstructural Analysis in NCD Detection
Individuals with NCDs exhibit speech deficits that impact spoken language across both acoustic and linguistic dimensions. Prior research has explored the various aspects of acoustic and linguistic features, encompassing prosody (e.g., dysfluencies and pauses), phonetics (e.g., Mel Frequency Cepstral Coefficients), phonology (e.g., mispronunciation of words) , voice quality (e.g., jitter and shimmer), lexicon (e.g., type-token-ratio and stopword ratio), syntax (e.g., parse-tree depth), and semantics (e.g., word and sentence embeddings) [32, 34, 43, 44, 12, 45, 46, 33, 47, 48, 41, 49, 50, 51]. These features at the microstructural level are essential for understanding the intricate language challenges faced by individuals with NCDs and serve as the foundation for microstructural analysis in this domain.
Recent efforts are devoted to identifying highly correlated and significant biomarkers of NCDs from the aforementioned features [52, 53, 54, 55]. For example, Weiner and Frankenberg et al. [53, 55] explored many traditional acoustic and linguistic features with a nested forward feature selection method, and found that the features on parts-of-speech, word categories and pauses are highly related to NCDs. While microstructural analysis provides valuable insights into specific acoustic and linguistic markers, it predominantly focuses on local features, and may overlook the broader narrative context.
@@ II-B Macrostructural Analysis in NCD Detection
While microstructural analysis highlights the specific language production impairments associated with NCDs, macrostructural analysis is equally important as it reveals how cognitive deficits impact higher-order cognitive processes such as attention, working memory, planning, organizing and structuring, which are crucial for constructing coherent, cohesive, and thematically consistent narratives [56, 57, 58, 31].
Macrostructural analysis in the literature generally focuses on features about cohesion (the use of linguistic devices for structural unity) and coherence (logical flow and overall consistency). Individuals with NCDs tend to produce text with impaired lexical cohesion (i.e., lexical repetition or substitution), wrong co-referencing, inconsistencies in maintaining the main topic, uncompleted or irrelevant information, and disrupted sequencing of events [59, 60, 61, 62, 63].
@@
Recent advances leverage machine learning and large language models (LLMs) for automatic macrostructure analysis on large corpora. The Coh-Metrix-Dementia program, with features like latent semantic analysis (LSA) and syntactic complexity, is well-established for extracting coherence metrics [27]. Some studies quantified global semantic coherence by using cosine similarity of word embeddings [38, 45], or evaluated pragmatic coherence via BLEU and METEOR scores using references [42]. More recent work adopted LLM as evaluator to automatically predict macro-descriptors [39].
However, as mentioned in Section I, these methods are unsuitable to NCD narratives due to frequent repetitions, unlimited references, and interpretability requirements. Additionally, there has been limited investigation into the temporal dynamics of topical evolution and the cross-modality alignment between visual stimuli and speech in VSN tasks. These gaps highlight the need for more advanced approaches to evaluate NCD narratives from both microstructural and temporal perspectives, addressing the interplay between evolving topics and the integration of visual and spoken elements.
@@ III Corpora
This study is conducted based on the English ADReSS (Alzheimer’s Dementia Recognition through Spontaneous Speech) Corpus [67], as well as the Cantonese CU-MARVEL-RABBIT (Cognitive Assessment Using Machine LeArning EmpoweRed VoicE AnaLysis – Rabbit Story Part) Corpus. ADReSS captures speech recorded from the Cookie Theft Picture Description Task [68]. CU-MARVEL captures narratives on the Rabbit Story – a storytelling task [21].
@@ III-A ADReSS
Figure 1: The Cookie Theft picture, where two kids are reaching for cookies, while a distracted mother is washing dishes, amidst a complex background.
@@
The ADReSS Corpus, developed by the University of Pittsburgh, contains transcripts of spontaneous speech samples from elderly English-speaking adults. The corpus is designed to facilitate research on age-related cognitive decline, focusing on NCDs (specifically, Alzheimer’s disease and mild cognitive impairment). The neurocognitive assessment tasks in this corpus involve the Cookie-Theft Picture Description (as shown in Fig. 1) for both HCs and NCDs groups. This corpus consists of audio recordings and transcriptions of interviews with 156 participants in the Cookie-Theft Picture Description task, balanced partitioned into training (54 HCs and 54 NCDs) and testing (24 HCs and 24 NCDs) subsets.
III-B CU-MARVEL’s Rabbit Story Task
The CU-MARVEL Corpus is designed and collected by our group in Hong Kong, with the participant inclusion criteria: (i) aged 60 or above; (ii) sufficient Cantonese language skills (listening and speaking), and (iii) sufficient vision and hearing ability (where glasses and hearing aids are acceptable) to complete all cognitive tests. This corpus contains series of cognitive tests, to align with the picture description task, we adopt the speech samples in the Rabbit Story task of CU-MARVEL, denoted as “CU-MARVEL-RABBIT” Corpus. It records Cantonese-speaking participants narrating a story about a girl finding a rabbit through 15 cartoon images, as illustrated in Fig. 2. These images are presented in a booklet, with each image displayed on a separate page. Participants will begin by reviewing all 15 images to familiarize themselves with the content. Subsequently, they will examine each image individually, providing a description of its content before proceeding to the next image. This process will continue sequentially until descriptions for all 15 images have been completed.
Figure 2: The Rabbit Story cartoon images – showing how a girl and her dog were trying to find a rabbit, and encountering various animals and adventures throughout their journey.
The CU-MARVEL-RABBIT Corpus differs from the ADReSS Corpus in its task design and analytical focus. While the ADReSS Corpus centers on single-image description to evaluate linguistic features associated with NCD, CU-MARVEL-Rabbit employs a multi-images description task, emphasizing the cognitive ability to connect sequential images and narrate a logical story. This task prioritizes macrostructural analysis, including coherence and cohesion, to assess narrative organization at a higher level. The version “2024-06-13” of the CU-MARVEL-RABBIT Corpus consists of 758 subjects (467 HCs, 291 NCDs), where the data distribution and the train-test split are listed in the Table I. The CU-MARVEL-RABBIT Corpus records the assessment for each subject based on specialists in neurology, along a scale from 0-4 regarding cognitive functioning, where the values 0 and 1 indicate HCs and 2, 3, and 4 indicate Neurocognitive Disorders (NCDs).

### v2

@@ Abstract
Early detection of neurocognitive disorders (NCDs) is crucial for timely intervention and disease management. Given that language impairments manifest early in NCD progression, visual-stimulated narrative (VSN)-based analysis offers a promising avenue for NCD detection. Current VSN-based NCD detection methods primarily focus on linguistic microstructures (e.g., pauses, lexical diversity), which are potentially linked to bottom-up (stimulus-driven) cognitive processing. While these features illuminate basic language abilities, the higher-order linguistic macrostructures (e.g., thematic or logical development), which may reflect top-down (concept-driven) cognitive abilities, remain underexplored. These patterns are crucial for NCD detection yet challenging to quantify due to their abstract and complex nature. To bridge this gap, we propose two novel dynamic macrostructural approaches: (1) Dynamic Topic Model (DTM) to track topic evolution over time, and (2) Text-Image Temporal Alignment Network (TITAN) to measure cross-modal consistency between speech and visual stimuli. Experimental results validated the efficiency of proposed approaches in NCD detection, with TITAN achieving superior performance both on the CU-MARVEL-RABBIT corpus (F1 = 0.7238) and the ADReSS corpus (F1 = 0.8889). The feature contribution analysis revealed that macrostructural features (e.g., topic variability, topic change rate, and topic consistency) constituted the most significant contributors in the model’s decision pathways, outperforming investigated microstructural features. These findings underscore the critical role of macrostructural patterns in understanding cognitive impairment mechanisms in NCDs.
@@ I Introduction
Neurocognitive disorders (NCDs)[1, 2], such as Alzheimer’s Disease, are highly prevalent and pose significant challenges to individuals and healthcare systems in our aging society. Early detection of these conditions is critical, as it can allow for timely intervention, potentially slowing disease progression and improving the quality of life for patients[3]. Traditional clinical approaches such as neuroimaging [4, 5], retinal photographs [6], electroencephalography (EEG) [7], cerebrospinal fluid (CSF) analysis [8], and in-person neuropsychological assessments [9] have been widely investigated. However, these approaches can be resource-intensive, invasive, and costly [10, 11]. Consequently, there is a growing emphasis on developing non-invasive, cost-effective, and scalable methods for early NCD detection.
One promising approach is the automatic analysis of speech and language, as communicative impairments are often among the earliest symptoms of cognitive decline [12]. These impairments may include temporal changes in spontaneous speech [13], word finding and retrieval difficulties [14], verbal dysfluencies [15], reduction in structural and thematic organization abilities [16, 17]. Such changes in speech patterns and language use provide valuable insights into underlying cognitive deficits and serve as potential biomarkers for early NCD detection. In light of this, various spoken cognitive tests, such as counting [18] and naming [19], have been designed to discover spoken language correlates for NCD detection. Among these, the visual-stimulated narrative (VSN) task [20, 21], such as describing a single image or telling a story based on image sequences, stands out by employing visual stimuli to elicit rich and spontaneous speech within a naturalistic and holistic framework. Specifically, the VSN task engages processes like viewing, describing, sequencing, and integrating elements, requiring multiple cognitive functions, such as attention, visual comprehension, language fluency, executive planning, memory, and social cognition [22, 23, 24, 25, 26, 27, 28]. Therefore, it offers superior potential for comprehensively evaluating cognitive abilities for NCD detection [29, 12].
Approaches to the analysis of VSNs can be categorized into two complementary aspects  [30]. First, the mainstream approach is microstructural analysis, which examines detailed, localized elements of a narrative, such as linguistic (e.g., lexicon, semantics, syntax) [31, 32] and acoustic (e.g., prosody, phonology) [33, 34] units. This level focuses on how things are produced and primarily reflects bottom-up (exogenous, stimulus-driven) cognitive processing. In contrast, macrostructural analysis evaluates the overarching organization of the narrative units, such as informativeness (overall information accuracy and completeness), coherence (logical flow of events or concepts) and text-visual consistency (alignment with visual stimuli [^1] This includes both depiction per image and connection across images, bridging informativeness and coherence.[35]. This level focuses on what is said and how things are combined, generally representing top-down (endogenous, concept-driven) cognitive integration.
Clinical research indicates that individuals with NCDs often exhibit greater impairment in macrostructuring than microstructuring. While microstructural deficits typically reveal specific language abilities (perception and production), macrostructural deficits more directly reflect impairments in higher-order cognitive abilities (attention, memory, organization, executive function), manifesting as disrupted topic maintenance, incoherent sequencing, and inconsistency to visual stimuli [30]. Despite this, macrostructural analysis remains underexplored, particularly within the context of NCD narratives.
@@
This research gap primarily arises from the inherent difficulty in quantifying macrostructural indicators due to their abstract and conceptual nature. Traditional approaches to quantify coherence rely on manual assessments, which are subjective and labor-intensive [36]. Recent advances in natural language processing (NLP) have introduced automated methods, such as calculating cosine similarities between consecutive textual embeddings or using large language models (LLMs) as evaluators [37, 38]. However, these methods often fall short for NCD narratives. Embedding-based similarities can misinterpret frequent repetitions or fillers as better cohesion and ignore dynamic evolutions of narrative topics over time. LLM-generated scores lack interpretability due to their untraceable decision pathways, undermining reliability in clinical applications.
@@
To quantify the consistency between narratives and visual stimuli, recent studies have used text-visual pretrained models to identify the highest relevance scores between image regions and sentences [39] or pattern-matching with fixed references [40, 41]. However, these methods have limitations: local text-image relevance fails to capture global consistency, and pattern-matching is restricted to fixed-reference tasks, making it unsuitable for creative tasks like storytelling. Therefore, tailored and automated approaches are needed for macrostructural analysis of NCD narratives.
@@
This work proposed to investigate specific cognitive and linguistic challenges by analyzing topical shifts, their temporal dynamics, and the coherence of narratives over time. We presented two novel approaches to macrostructural narrative analysis: Dynamic Topic Models (DTM) for temporal analysis of topical evolution; and Text-Image Temporal Alignment Model (TITAN) for cross-modal (i.e., stimulus images and spoken narratives) alignment analysis. This approach aims to reveal cognitive deficits by identifying narrative impairments associated with NCDs, and to provide a deeper understanding of their impact on communication and cognition. The investigations are conducted using the Rabbit Story task (a VSN task) from the CU-MARVEL corpus (a Cantonese corpus), as well as the single-image description task in the ADReSS corpus (an English corpus). Our experiments validated the effectiveness of dynamic topic consistency as a macrostructure metric (F1=0.65, AUC=0.78). The proposed TITAN model achieved the highest performance (F1=0.72, AUC=0.81) in classification tasks, surpassing those of established microstructure and handcrafted macrostructure feature sets. Moreover, experiments of cross-comparison and regression tasks further demonstrated the effectiveness of the proposed dynamic macrostructure modeling approaches for NCD detection.
@@
The paper is organized as follows: Section II reviews related works, Section III introduces corpora used in this study, Section IV introduces the proposed approaches, Section V presents our experimental settings and results, Section VI analyzes and discusses experimental results, and Section VII concludes this work and prospects future works.
@@ II Related Work
Extensive research has been conducted for NCD detection based on spoken language, leveraging various methodologies to identify biomarkers and assess cognitive impairments. This section reviews pertinent literature encompassing both microstructural and macrostructural narrative analysis, collectively laying the groundwork for enhanced narrative analyses. The term “microstructure” refers to fine-grained elements of narrative production, including lexical, syntax, semantics, phonetics, acoustics, and other smaller linguistic units. In contrast, the term “macrostructure” refers to higher-level aspects of narrative construction, such as coherence, cohesion, thematic development, and the overall organization of narratives. Both microstructural and macrostructural analysis are crucial for a comprehensive understanding of cognitive deficits reflected in NCD narratives. However, existing studies often face limitations in capturing the full spectrum of narrative features, particularly the dynamic aspects of narrative construction. By identifying these limitations, this review highlights the gaps the proposed methods aim to fill, specifically by incorporating temporal dynamics and macrostructural consistency, thus advancing the narrative analysis of NCDs.
@@ II-A Microstructural Analysis in NCD Detection
Individuals with NCDs exhibit speech deficits that impact spoken language across both acoustic and linguistic dimensions. Prior research has explored the various aspects of acoustic and linguistic impairments in NCD speech, encompassing prosody (e.g., increased dysfluencies and longer, more frequent pauses), phonetics (e.g., abnormal Mel Frequency Cepstral Coefficients patterns), phonology (e.g., higher rates of mispronunciation of words), lexicon (e.g., reduced type-token-ratio and elevated stopword ratio), syntax (e.g., reduced parse-tree depth), and semantics (e.g., abnormal patterns in word and sentence embeddings) [31, 42, 33, 43, 44, 13, 45, 46, 32, 47, 48, 40, 49, 50, 51]. These microstructural features are essential for understanding the language abilities of individuals with NCDs, providing the foundation for analysis in this domain.
Recent efforts are devoted to identifying highly correlated and significant biomarkers of NCDs from the aforementioned features [52, 53, 54, 55, 56]. For example, Cho et al. [56] linked higher CSF p-tau [^2] CSF p-tau: Cerebrospinal fluid phosphorylated tau, a NCD-related biomarker indicating the burden of neurofibrillary tangle pathology. levels (signifying greater burden of NCD pathology) to specific acoustic and linguistic features: lower preposition and noun ratios, higher pause rates, and shorter speech segment durations. Such microstructural analysis provides valuable insights into local features, but more comprehensive information may be offered by macrostructural approaches that examine broader narrative structure and contexts. Hence, this work extends prior research by examining VSN’s macrostructure such as topic evolution and cross-modal consistency.
(a) The Cookie Theft picture
(b) The Rabbit Story cartoon
Figure 1: Visual stimuli used for NCD detection: (a) Cookie Theft (depicting two kids stealing cookies while a distracted mother is washing dishes, amidst a complex background) and (b) Rabbit Story (showing a girl and her dog searching for a rabbit and encountering various animals and adventures).
@@ II-B Macrostructural Analysis in NCD Detection
While microstructural analysis highlights specific language production impairments associated with NCDs, macrostructural analysis is equally important as it reveals how cognitive deficits affect higher-order cognitive processes like attention, working memory, planning, and organization. This macrostructural perspective is essential for understanding how NCDs affect the construction of coherent, cohesive, and thematically consistent narratives [57, 58, 59, 30].
Macrostructural analysis in the literature generally focuses on features about cohesion (the use of linguistic devices for structural unity) and coherence (logical flow and overall consistency). Individuals with NCDs tend to produce text with wrong co-referencing, inconsistencies in maintaining the main topic, uncompleted or irrelevant information, and disrupted sequencing of events [60, 61, 62, 63, 64].
@@
Recent advances leverage machine learning and large language models (LLMs) for automatic analyze macrostructure of VSNs. Some studies have quantified global semantic coherence by using cosine similarity of consecutive embeddings [37, 45], while others have assessed pragmatic coherence via BLEU and METEOR scores using references [41]. More recent work has adopted LLM as evaluator to automatically predict macro-descriptors [38].
However, as mentioned in Section I, these methods are unsuitable for NCD spoken narratives: similarity scores are confounded by frequent repetitions, reference-based metrics (BLEU, METEOR) are inapplicable due to non-standardized references, and LLM-based evaluation fails to meet interpretability requirements. Additionally, there has been limited investigation into the temporal dynamics of topical evolution and the cross-modality alignment between visual stimuli and speech in VSN tasks. These gaps highlight the need for more advanced approaches to evaluate NCD narratives from both macrostructural and temporal perspectives, addressing the interplay between evolving topics and the integration of visual and spoken elements.
@@ III Corpora
This study is conducted based on the ADReSS (Alzheimer’s Dementia Recognition through Spontaneous Speech) Corpus [68], as well as the CU-MARVEL-RABBIT (Cognitive Assessment Using Machine LeArning EmpoweRed VoicE AnaLysis – RABBIT Story Part) Corpus. These corpora were selected to enable a comprehensive understanding of VSN patterns for NCDs across different language and narrative types. ADReSS collects English recordings from the Cookie Theft picture description task [69], while CU-MARVEL collects Cantonese narratives from the Rabbit Storytelling task [70].
@@ III-A ADReSS
The ADReSS Corpus, developed by the University of Pittsburgh, contains transcripts of spontaneous speech samples from elderly English-speaking adults. The corpus is designed to facilitate research on age-related cognitive decline, focusing on NCDs (specifically, Alzheimer’s disease and mild cognitive impairment). The neurocognitive assessment tasks in this corpus involve the Cookie-Theft Picture Description, as shown in Fig. 1(a), for both Healthy Control (HC) and NCD groups. This task assesses participants’ ability to recognize the elements and produce a coherent narrative based on their spatial relationships and implied events. This corpus consists of audio recordings and transcriptions of interviews with 156 participants in the Cookie-Theft Picture Description task, balanced partitioned into training (54 HCs and 54 NCDs) and testing (24 HCs and 24 NCDs) subsets.
III-B CU-MARVEL-RABBIT
The CU-MARVEL Corpus is designed and collected by our group in Hong Kong, encompassing various cognitive tests. To assess the cognitive capacities of individuals with NCDs in integrating sequential visual information and constructing coherent narratives, we adopt the Rabbit Story task of CU-MARVEL, denoted as “CU-MARVEL-RABBIT” Corpus.
The corpus records Cantonese-speaking participants narrating a story about a girl finding a rabbit through 15 cartoon images, as illustrated in Fig. 1(b). These images generally depict 5 sequential scenes (or topics): (1) the discovery and subsequent loss of the rabbit (images 1-4), (2) an encounter with a squirrel and some ants (images 5-7), (3) a meeting with a goose (images 7-9), (4) a run-in with a cow (9-11), and (5) a reunion with the rabbit followed by a farewell wave to its family (images 12-15).
These images are presented in a booklet, with each image displayed on a separate page. Participants will first review all 15 images to familiarize themselves with the content. Then, they will examine each image individually, providing a description of its content before proceeding to the next image. This process will continue sequentially until the stories for all 15 images have been completed.
Figure 2: Illustrative Cantonese narratives in CU-MARVEL-RABBIT, where English words are roughly translated by GPT-4o for reader’s comprehension only.
The CU-MARVEL-RABBIT Corpus (version “2024-06-13”) consists of 758 subjects, including 467 HCs and 291 individuals with NCDs. Illustrative examples in Fig. 2, where the English translations are coarsely rendered by GPT-4o [71] for reader’s comprehension only, highlight the typical differences between HC and NCD narratives. Specifically, HC tends to capture most visual details (e.g., characters (girl, dog, rabbit, squirrel, goose, cow), objects (tree hollow, rock), and settings (house, garden, forest)) and weaves them into a clear, logical storyline across the images (e.g., rabbit was missing → girl and dog searched in the forest → rabbit was found with its family). In contrast, NCD often misses or misidentifies key details (e.g., squirrel, goose) and exhibits difficulties in establishing coherent connections between events (primarily isolated descriptions).
The data distribution and train-test split are listed in Table I. To benchmark the performance of frontend tools (including speech recognition, word segmentation, and part-of-speech tagging), we extensively annotated a subset of 46 participants (22 HCs, 24 NCDs) from the test set (aligned with [48]), naming “Frontend-Evaluation” set. The CU-MARVEL-RABBIT Corpus records the assessment for each subject by specialists in neurology, using a 0–4 severity scale for cognitive impairment, with lower values corresponding to better cognitive performance. Specifically, values 0-1 indicate HCs and 2, 3, and 4 represent progressively severe NCDs.